OpenStudy (anonymous):

Identify the vertex for the graph of y = −3x2 + 6x − 4.

OpenStudy (anonymous):

magic formula for vertex is \(-\frac{b}{2a}\) for the first coordinate second coordinate is what you get when you repalace \(x\) by \(-\frac{b}{2a}\)

OpenStudy (anonymous):

To find the x cord of the vertex use \[ -\frac{b}{2a} \] Just plugin your b value and a value \[ -\frac{6}{2(-3)} = 1\] \[ x = 1\] Now put the value of x, which is 1 in to the x variable of the equation for find y \[ y = −3x^2 + 6x − 4 \] \[ y = −3(1)^2 + 6(1) − 4 \] [ y = -1 \] So our vertex is at (1,-1) Also the parabola will open down since a(-3) is a negative and we will have a maximum.